Solve for the equation of the orthogonal trajectory: 1. Y = 3x - 1 + Ce-3x
1
Expert's answer
2020-12-02T01:46:57-0500

Given: "y=3x-1+Ce^{-3x}" ... (1)

Find the differential equation for the given family of curves by eliminating the arbitrary constant "C" .

Differing the given equation (1) with respect to "x" , we get

"\frac{dy}{dx}=3(1)-0+Ce^{-3x}(-3)=3-3Ce^{-3x}"

That is "\frac{dy}{dx}=3-3Ce^{-3x} ... (2)"

Eliminate "C" from the equations (1) and (2).

From equation (2), "3Ce^{-3x} =3-\frac{dy}{dx} \Rightarrow C=e^{3x}-\frac{1}{3}e^{3x}\frac{dy}{dx}"

Substitute "C" in equation (1), we get "y=3x-1+\left ( e^{3x}-\frac{1}{3}e^{3x}\frac{dy}{dx} \right )e^{-3x}=3x-1+1-\frac{1}{3}\frac{dy}{dx}"

"\Rightarrow y=3x-\frac{1}{3}\frac{dy}{dx}"

Replace "\frac{dy}{dx}" by "-\frac{dx}{dy}" to get the differential equation of the orthogonal trajectories.

Then, we get

"y=3x-\frac{1}{3}\left ( -\frac{dx}{dy} \right )=3x+\frac{1}{3}\frac{dx}{dy}"

"\frac{dx}{dy}+9x=3y"

The above is a first order linear differential equation of the form "\frac{dx}{dy}+p(y)x=g(y)" ,

where "p(y)=9" and "g(y)=3y"

General solution to the linear differential equation is

"x(" Integrating Factor) = "\int (" Integrating Factor)("g(y))dy+c" ,

where Integrating Factor = "e^{\int p(y)dy}"

Using the above, Integrating Factor = "e^{9y}"

General solution is "x(e^{9y})=\int (e^{9y})(3y)dy+c"

Use integration by parts to evaluate the right side integral.

Then, we get

"x=\frac{y}{3}-\frac{1}{27}+ce^{-9y}"

Therefore, orthogonal trajectories for the given family of curves is

"x=\frac{y}{3}-\frac{1}{27}+ce^{-9y}" , where "c" is arbitrary constant.